# Apparent teleportation of indistinguishable particles

### v2

@@ Abstract
Teleportation, introduced in science fiction literature, is an instantaneous change of the position of a macroscopic object. Two teleportation-like phenomena were predicted by quantum mechanics: quantum teleportation and, recently, quantum particle teleportation. The former is investigated experimentally and has applications in quantum communication and computing.
@@
Here, we introduced the third teleportation-like phenomenon—apparent teleportation. It seems to be a natural consequence of the Standard Model’s indistinguishable elementary particles and antiparticles. We give an example of a process leading to the apparent teleportation within a 1+1D toy model of particle space-time evolution. It utilizes the local transport of particles and antiparticles and the local creation and annihilation of particle-antiparticle pairs. Furthermore, we suggest a method to observe the apparent teleportation in nucleus-nucleus collisions at properly selected collision energies. The method requires the measurement of correlations between momenta of charm and anticharm hadrons in collisions with a single $c\bar{c}$ pair being produced. The ultimate prediction following the apparent teleportation hypothesis is the uncorrelated emission of charm and anticharm hadrons. It can be tested by contemporary experiments.
Observing the apparent teleportation would uncover the basic transport properties of indistinguishable particles. In particular, the apparent teleportation may explain the rapid thermalisation of the system created in collisions of two atomic nuclei and the quark-gluon plasma hadronisation paradox. Theoretical and experimental efforts are needed to observe the apparent teleportation processes and study their properties.
@@
Teleportation [1] is rooted in science fiction. It describes the ability of macroscopic objects, for example, humans, to reappear at some distant location without traversing space. Thus, the speed of light does not restrict the possibility of changing our location in space. Needless to say, this also concerns signals. Details depend on the science fiction setup.
@@
Quantum teleportation [2] is the standard quantum physics phenomenon - transferring quantum-state properties over arbitrarily large distances in no time. In quantum teleportation, conditional probabilities of measurement outcomes reappear at a distant location using the entanglement of two particles. There is neither a super-luminal transfer of conserved quantities nor super-luminal signalling. The quantum teleportation is measured in numerous experiments [3, 4, 5, 6].
@@
Quantum particle teleportation [7] is an instantaneous transfer of a single particle over a large distance predicted by standard quantum mechanics. The phenomenon appears for particles in a potential. It is induced by frequently monitoring if the particle is at rest (a peculiar quantum Zeno dynamics [8, 9, 10]). Quantum particle teleportation has the properties of ”science fiction” teleportation, allowing for the reappearance of conserved quantities and signalling. It was predicted in 2023 and has not yet been tested experimentally.
@@
Apparent teleportation (AT) is a concept introduced here. It is defined as the time evolution of a system of indistinguishable particles that cannot be mimicked by distinguishable particles moving with luminal or subluminal velocities. Of course, when referring to distinguishable particles mimicking indistinguishable ones, we assume the observer is blind to particle labels. The apparent teleportation seemingly requires superluminal velocities of particles. This illusion results from particle trajectories and velocities being defined for distinguishable particles but undefined for indistinguishable ones.
I Results
@@
Elementary particles and antiparticles of the Standard Model are indistinguishable bosons and fermions. Objects composed of many indistinguishable elementary particles and antiparticles are close-to-unique and thus can be approximately treated as distinguishable ones 111 An internally consistent model covering distinguishable and indistinguishable particles and antiparticles has to give identical predictions for a single particle or antiparticle in the system when treating it as distinguishable and indistinguishable.. Classical physics is rooted in our daily experience within the world of macroscopic distinguishable humans and other objects. In particular, it is well established that they can move only with velocities smaller than the speed of light. However, applying time evolution limits of distinguishable objects to indistinguishable particles and antiparticles imposes human-centered experience on a fundamentally different quantum world. In particular, the problem is obvious if these limits are given in quantities undefined for indistinguishable particles and antiparticles, like particle velocity. In this sense, the apparent teleportation is the natural property of models postulating the indistinguishability of elementary particles and antiparticles.
@@
In the following, we give a simple example of a model predicting the apparent teleportation. It is a toy model of particle-antiparticle creation, annihilation, and transport. Secondly, we present an idea of experimentally observing the apparent teleportation in collisions of two nuclei.
I.1 Toy model of apparent teleportation
@@
Cell Model is a dynamical model based on the 1+1D discrete-time Markov chain framework [12, 13]. Space is assumed to be a vector of $V$ discrete cells $(v_{1},v_{2},\ldots,v_{V})$ arranged in a ring. At the given time $t$, the system microstate is fully defined by a distribution of particles and antiparticles in cells. The system’s evolution in time is assumed to be discrete, and the time steps are numbered by $t$. During evolution, transitions between the systems’ microstates occur. The transition probability from a microstate $X$ at $t$ to a microstate $Y$ at $t+1$ depends only on the microstate $X$ the basic assumption of Markov’s chains.
@@
For this paper, we consider three types of reactions changing the system’s microstates:
@@
The first two reactions change the number of particles and antiparticles, whereas the third only changes their distribution in cells. The difference between particle and antiparticle numbers of a given type is conserved in the whole system. This is the only conservation law imposed on the system. In relation to quantum mechanics, the time steps are considered times at which a measurement of (anti-)particles’ positions in space is performed, and the uncertainty principle gives the size of cells in space.
In summary, the distribution of particles and antiparticles in cells differs at different time steps. These changes can be due to particle-antiparticle creation-annihilation reactions and particle and antiparticle redistribution. Now, we postulate that only changes obeying the transport locality are allowed during the redistribution. For distinguishable particles, the transport-locality requirement reduces to a requirement of each particle moving by no more than $\Delta$ cells. In physics, this corresponds to particle velocities, which are limited by the speed of light in the vacuum.
For indistinguishable particles, the particle’s trajectories and velocities are undefined. Thus, the transport locality condition for distinguishable particles does not apply to indistinguishable ones. The condition for indistinguishable particles and antiparticles is unknown, and experiments are needed to uncover it.
Here, we assume that the transport locality condition introduced for a conserved charge without considering particle-antiparticle creation and annihilation processes [13] (see below) is valid separately for redistributions of indistinguishable particles and antiparticles. It implies the following. During a single time step, the particle (antiparticle) number in any interval of cells cannot be transported beyond an interval by $\Delta$ cells longer on the left and right. Moreover, it cannot be squeezed to an interval by $\Delta$ cells shorter on the left and right. The condition means that only those redistributions of particles (antiparticles) that can be mimicked by distinguishable particles (antiparticles) are possible. The condition is given by two transport-locality inequalities [13]:
$\sum_{j=i}^{i+k}n_{j}^{X}\leq\sum_{l=i-\Delta}^{i+k+\Delta}n_{l}^{Y}\leavevmode\nobreak\ ,$ (1)
$\sum_{l=i}^{i+k}n_{l}^{Y}\leq\sum_{j=i-\Delta}^{i+k+\Delta}n_{j}^{X}\leavevmode\nobreak\ ,$
@@
Figure 1: The simple setup illustrates the apparent teleportation of indistinguishable particles within the Cell Model. The maximum distance a distinguishable particle and antiparticle can move during a time step is set to $\Delta=1$. Three types of transitions, pair creation, local transport and pair annihilation, are possible. Left: The example process with indistinguishable particles starting from a single particle in cell three at $t=0$ and ending in a single particle in cell seven at $t=3$. Right: Unsuccessful attempt to mimic the process by distinguishable particles. Three pairs of particles and antiparticles with identical labels appear at $t=1$. Particles are shifted by one cell to the right at $t=2$. In particular, the particle ’d’ appears in cell seven. The particle number distributions for indistinguishable and distinguishable particles are identical at $t=0$, 1 and 2. The pair annihilation reactions in cells 4, 5 and 6 are impossible during the third step because cells contain particles and antiparticles with different labels. For example, cell 4 contains particle ’a’ and antiparticle ’b’ at $t=2$. See the text for more details.
@@
Apparent teleportation within Cell Model is illustrated in Fig. 1. An example of the system’s time evolution with $V=9$ cells and periodic boundary conditions is shown. Cells contain ”background” particles carrying total energy depicted by the grey band. Particles and antiparticles of interest (for example, $c$ and $\bar{c}$ quarks, see below) are depicted by red and blue circles, respectively. For simplicity, we assume that the speed of light results in $\Delta=1$. The difference between the total number of particles and antiparticles is assumed to be conserved.
@@
The case of indistinguishable particles and antiparticles is presented in the left plot of Fig. 1. We postulate that only one particle exists at $t=0$, and it is in cell 3. Three pairs of particles and antiparticles appear in cells 4, 5 and 6 at $t=1$. At $t=2$, the particle number distribution is shifted by one cell to the right by the transport-local transition with $\Delta=1$. This is allowed by the transport-locality inequalities (1). Then, at $t=3$, particle-antiparticle pairs annihilate, leaving a single particle in cell seven.
@@
The right plot of Fig. 1 shows an unsuccessful attempt to mimic the above process using distinguishable particles. In the example presented here, they are labeled by the letters ’a’, ’b’, ’c’, and ’d’. Similarly, antiparticles have unique labels. Only creating a particle and an antiparticle with identical labels is possible. Similarly, only a particle and an antiparticle with identical labels can annihilate.
@@
For the label-blind observer, the particle and antiparticle number distributions of indistinguishable and distinguishable particles are identical at $t=0,1$, and 2. But the similarity breaks at $t=3$. This is because annihilating particles and antiparticles with different labels is impossible. Of course, distinguishable particles can mimic the process depicted in the left plot if the transport-locality requirement is lifted. Particle ’a’ can be ’teleported’ by four cells in three-step steps from cell $3$ to $7$.
@@
The example construction of the process shown in Fig. 1 can be easily generalized, allowing for an arbitrary redistribution of indistinguishable boson-like particles and antiparticles in three-time steps. This includes apparent teleportations - the processes which seemingly violate the speed of light limit. Their rate will increase with the increasing rate of creation and annihilation reactions and thus with energy density. Following the quantum mechanics, the creation and annihilation of virtual (disobeying conservation laws for a short time) particles and antiparticles is also possible.
@@
We note that the microscopic mechanism of the apparent teleportation illustrated in Fig. 5 (left) cannot be used for superluminal signalling. An observer in cell seven detecting a particle at $t=3$ cannot be sure it is caused by injecting a particle at $t=0$ in cell three. For example, it could be due to the creation of particle and antiparticle pair in cell six independently of events in cell three and then a particle appearing in cell seven.
@@
The probability of a given process depends on the dynamics. One can speculate that at sufficiently high energy densities, the system of indistinguishable particles approaches maximum symmetry (the microstate symmetry [13]), and all processes have the same probability. This would imply that the system is ”born in equilibrium” [14] - the probability of any microstate appearing at any time is equal. The apparent teleportation also seem to explain the quark-gluon paradox formulated in Ref. [15]. The transition of large volume quark-gluon plasma created in collisions of two atomic nuclei (see below for details) to Colorless hadrons require ”teleportation” of colour charge to assure local colour neutrality. Quarks and gluons are indistinguishable particles. Thus, the apparent teleportation can serve as the physical mechanism of the ”teleportation”.
The above discussion suggests that the experimental search for apparent teleportation is the most promising at high energy densities, where the particle and antiparticle density is high, and their creation and annihilation reactions are frequent.
I.2 Observing apparent teleportation
@@
System of quarks and gluons created in nucleus-nucleus collisions at high energies is the highest energy density system ($\epsilon>1$ GeV/fm3 $\approx 1.6\cdot 10^{35}$ J/m3) created under conditions controlled in the laboratory [16]. Its equilibrium state is called quark-gluon plasma (QGP) [17]. Its total charge of strong interactions, colour, is zero. The plasma expands and cools down, and the transition to colour-neutral hadrons - small bags of quarks and gluons -occurs at the hadronisation (transition) temperature $T_{H}\approx 150$ MeV [18].
@@
Six types of quarks (fermions) $u$, $d$, $s$, $c$, $t$ and $b$ and the corresponding antiquarks and the eight types of gluons (bosons) can be created during the collision. Particles and antiparticles of a given type are indistinguishable. Massless gluons and light $u$ and $d$ quarks and antiquarks are the most abundant. Production of other quarks is suppressed even at the top collision energies of the CERN LHC [19]. The threshold for the QGP creation in heavy-ion collisions is located at $\sqrt{s_{NN}}\approx 10$ $\mbox{Ge\kern-1.00006ptV}\!/\!c$ [20, 21], for review see also Refs. [22, 23]. The fixed target experiments at the CERN SPS measure heavy-ion collisions in the energy range $\sqrt{s_{NN}}\approx 5-20$ $\mbox{Ge\kern-1.00006ptV}\!/\!c$. At the top SPS energy, just above the QGP threshold energy, the mean number of light quarks and gluons produced in central collisions of two lead nuclei is on the order of 1000 [24]. The corresponding numbers for strange and charm quarks are $\approx$ 100 [24] and $\approx$ 1 [25], respectively.
@@
The local creation and annihilation of $c$ and $\bar{c}$ pairs mean that the energy and momentum scale of the interaction determines $c$-$\bar{c}$ space-time separation via the uncertainty principle. The scale ranges between:
@@
    The theoretical challenge: Quantum Chromodynamics, the theory of strong interactions, does not provide quantitative predictions for multi-particle correlations in heavy-ion collisions; for further discussion, see Ref. [26].
However, with the help of well-established heavy-ion models, it may be possible to provide experimental evidence of the apparent teleportation by measuring momentum correlations between $c$ and $\bar{c}$ quarks. The idea is rooted in our previous paper [26], which suggested the use of momentum correlations of charm-anticharm hadrons to extract information on spatial correlations of $c$ and $\bar{c}$ quarks at hadronisation. The paper reviews the experimental results on charm hadron correlations and collective flow. It also discusses the importance of having collisions with no more than one $c\bar{c}$ pair created. We note that the requirement of a single $c\bar{c}$ pair in a collision can be reconciled with the example of the apparent teleportation mechanism presented in Sec. I.1, assuming that virtual $c\bar{c}$ pairs are produced at a sufficiently high rate. Below, we detail the idea.
Figure 2: The simplest illustration of observing the apparent teleportation in heavy-ion collisions at high energies using a single pair of $c$ and $\bar{c}$ quarks. The top plot shows the space-time (1+1D) evolution of the quark-gluon plasma created at $t=0$ in the spatial region from $-R$ to $R$. The hadronisation is assumed to occur at $t=R$ when the system extends from $-2R$ to $2R$. The middle plot shows an example of the positions of the $c$ and $\bar{c}$ quarks at hadronisation, denoted as $r_{c}$ and $r_{\bar{c}}$, respectively. In this example, the past light cones of the quarks overlap within the collision light cone. This overlap is required for the subluminal movement of distinguishable $c$ and $\bar{c}$ quarks created at a common space-time point. We refer to this as the local indistinguishability condition (LD condition). The bottom plot depicts the positions of the $c$ and $\bar{c}$ quarks, which can only be explained by the apparent teleportation.
@@
    Initial stage: a high-density quark-gluon plasma is created.
@@
    Expansion stage: the plasma expands [27], cooling down to the hadronisation temperature ($T_{H}\approx 150$ MeV in the local frame of the flowing matter). This stage is modelled using relativistic hydrodynamics [16].
@@
    Hadronisation stage: the plasma is converted into hadrons and resonances following statistical rules [28, 29] applied in the rest frame of a plasma fluid element. As a result, hadron momenta are determined by their flow acquired during the expansion stage and by the hadronisation (a local statistical process).
@@
We illustrate the idea of observing the apparent teleportation with the help of the schematic 1+1D model of the collisions sketched in Fig. 2. For simplicity, we assume that the initial stage is created at $t=0$ and extends in space from $-R$ to $R$. The hadronisation occurs at $t=R$; at this stage, the system extends from $-2R$ to $2R$. For simplicity, we neglect the finite size of the space-time cell of the local $c\bar{c}$ pair creation. In the middle and bottom plots, two examples of the hadronisation points of a single $c\bar{c}$ pair created in the collision are depicted as $r_{c}$ and $r_{\bar{c}}$. The plots also show the past light cones of the charm and anticharm quarks (the blue and red areas).
Let us assume the $c$ and $\bar{c}$ quarks were created in the same space-time point within the collision light cone and moved to hadronisation as distinguishable particles with subluminal velocities. In this case, the creation point has to be located inside the overlap between the collision light cone and the past-light cones of $c$ and $\bar{c}$ quarks - the local-distinguishable (LD) condition. The $c\bar{c}$ pair depicted in the middle plot of Fig. 2 fulfils the requirement, but the pair depicted in the bottom plot does not. The latter implies that the process was not local-distinguishable, and one needs the apparent teleportation to explain it.

### v4

# Kielce, Poland

@@ Abstract
Teleportation, introduced in science fiction literature, is an instantaneous change of the position of a macroscopic object. Two teleportation-like phenomena have been predicted by quantum mechanics: quantum teleportation and, more recently, quantum particle teleportation.
@@
Here, we introduce the third teleportation-like phenomenon—apparent teleportation. It seems to be a natural consequence of the Standard Model’s indistinguishable elementary particles and antiparticles. We illustrate the idea within a 1+1D toy model of particle-antiparticle creation and space-time evolution obeying transport locality. Furthermore, we propose a novel method to observe apparent teleportation driven by strong interactions through measurements of correlations between the momenta of charm and anticharm hadrons in nuclear collisions. Observing the apparent teleportation would uncover the basic transport properties of indistinguishable particles.
arxiv: 2503.10565
1 Introduction
@@
Teleportation Kaku2008 is rooted in science fiction. It describes the ability of macroscopic objects, such as humans, to reappear at a distant location without traversing space. Thus, the speed of light does not restrict the possibility of changing our location in space. Needless to say, this also concerns signals. Details depend on the science fiction setup.
@@
Quantum teleportation Bennett:1992tv is the standard quantum physics phenomenon - transferring quantum-state properties over arbitrarily large distances in no time. In quantum teleportation, conditional probabilities of measurement outcomes reappear at a distant location using the entanglement of two particles. There is neither a super-luminal transfer of conserved quantities nor super-luminal signalling. The quantum teleportation is measured in numerous experiments Bouwmeester:1997slj ; Yin:2017ips ; Pfaff:2014lgp ; Pirandola:2015nwo .
@@
Quantum particle teleportation Porras:2023fgn is an instantaneous transfer of a single particle over a large distance predicted by standard quantum mechanics. The phenomenon appears for particles in a potential. It is induced by frequently monitoring if the particle is at rest (a peculiar quantum Zeno dynamics Facchi:2000bs ; Facchi:2002jgu ; Facchi:2008nrb ). Quantum particle teleportation exhibits the properties of ”science fiction” teleportation, allowing for the reappearance of conserved quantities and signalling. It was predicted in 2023 and has yet to be experimentally tested.
@@
Apparent teleportation (AT) is a concept introduced here. It is defined as the time evolution of a system of indistinguishable particles that cannot be mimicked by distinguishable particles moving with luminal or subluminal velocities. Of course, when referring to distinguishable particles mimicking indistinguishable ones, we assume the observer is blind to particle labels. The apparent teleportation seemingly requires superluminal velocities of particles. This illusion arises from the fact that particle trajectories and velocities are defined for distinguishable particles but are undefined for indistinguishable ones.
2 Results
@@
Elementary particles and antiparticles of the Standard Model are indistinguishable bosons and fermions. Objects composed of many indistinguishable elementary particles and antiparticles are close to unique and thus can be approximately treated as distinguishable ones 111 An internally consistent model covering distinguishable and indistinguishable particles and antiparticles has to give identical predictions for a single particle or antiparticle in the system when treating it as distinguishable and indistinguishable.. Classical physics is rooted in our daily experience within the world of macroscopic, distinguishable objects, including humans. In particular, it is well established that they can move only with velocities smaller than the speed of light. However, applying time evolution limits of distinguishable objects to indistinguishable particles and antiparticles imposes human-centred experience on a fundamentally different world. The inadequacy is obvious if these limits are given in quantities that are undefined for indistinguishable particles and antiparticles, such as particle velocity. In this sense, apparent teleportation is the natural property of models postulating the indistinguishability of elementary particles and antiparticles.
@@
In the following, we illustrate the concept of apparent teleportation within a toy model of particle-antiparticle creation, annihilation, and transport. Importantly, we propose a novel method which allows for observing the apparent teleportation in collisions of two nuclei, if it exists.
2.1 Toy model of apparent teleportation
@@
Cell Model is a dynamical model based on the 1+1D discrete-time Markov chain framework Gazdzicki:2017rfe ; Gazdzicki:2022zej . Space is assumed to be a vector of $V$ discrete cells $(v_{1},v_{2},\ldots,v_{V})$ arranged in a ring. At the given time $t$, the system microstate is fully defined by a cell distribution of particles and antiparticles. The system’s evolution in time is assumed to be discrete, and the time steps are numbered by $t$. During evolution, transitions occur between microstates. The transition probability from a microstate $X$ at $t$ to a microstate $Y$ at $t+1$ depends only on the microstate $X$ the basic assumption of Markov chains.
@@
To illustrate the concept of apparent teleportation, we consider three types of reactions changing the system’s microstates:
@@
The first two reactions change the number of particles and antiparticles, whereas the third only changes the distribution of (anti) particles in cells. The difference between particle and antiparticle numbers of a given type is conserved in the whole system. This is the only conservation law that applies to the system.
To summarise, the distribution of particles and antiparticles in cells generally differs at different time steps. This can be due to particle-antiparticle creation-annihilation reactions and particle and antiparticle redistribution. Now, we postulate that only changes obeying the transport locality are allowed during the redistribution. For distinguishable particles, the transport-locality requirement reduces to a requirement of each particle moving by no more than $\Delta$ cells. In physics, this corresponds to particle velocities, which are limited by the speed of light in a vacuum.
For indistinguishable particles, the particles’ trajectories and thus velocities are undefined. Thus, the transport locality condition for distinguishable particles does not apply to indistinguishable ones. The condition for indistinguishable particles and antiparticles is unknown, and experiments are needed to uncover it.
Here, we assume that the transport locality condition introduced for conserved particle number in Ref. Gazdzicki:2022zej (see below) is valid separately for redistributions of indistinguishable particles and antiparticles. It implies the following. During a single time step, the particle (antiparticle) number in any interval of cells cannot be transported beyond an interval by $\Delta$ cells longer on the left and right. Moreover, it cannot be squeezed into an interval by $\Delta$ cells shorter on the left and right. The condition means that only those redistributions of particles (antiparticles) that can be mimicked by distinguishable particles (antiparticles) are possible. The condition is given by two transport-locality inequalities Gazdzicki:2022zej :
$\sum_{j=i}^{i+k}n_{j}^{X}\leq\sum_{l=i-\Delta}^{i+k+\Delta}n_{l}^{Y}~,$
$\sum_{l=i}^{i+k}n_{l}^{Y}\leq\sum_{j=i-\Delta}^{i+k+\Delta}n_{j}^{X}~,$ (1)
@@
Concerning quantum mechanics, the Cell Model for indistinguishable particles mimics the time evolution of a coherent quantum state of particles on a one-dimensional space lattice. Not considering particles’ momenta and initially allowing for all transitions corresponds to having the wave functions of (anti)particles delocalise in the whole system. The Heisenberg uncertainty principle provides a simple and intuitive lower bound on momentum, which relates directly to quantum de-localisation and coherence: $\Delta p\approx\hbar/(2L)$, where $L$ is the system size. The time steps correspond to the time intervals between measurements of (anti) particle positions. Out of all possible transitions, the transport-locality condition (2.1) allows only those that obey the speed-of-light limit. This, together with the particle-antiparticle creation and annihilation processes, provides the model with the basic properties of quantum mechanics and relativistic physics.
@@
Figure 1: The simple setup illustrates the apparent teleportation of indistinguishable particles within the Cell Model. The maximum distance a distinguishable particle and antiparticle can move during a time step is set to $\Delta=1$. Three types of transitions — pair creation, local transport, and pair annihilation — are possible. a The example process with indistinguishable particles starting from a single particle in cell three at $t=0$ and ending in a single particle in cell seven at $t=3$. b Unsuccessful attempt to mimic the process by distinguishable particles. Three pairs of particles and antiparticles with identical labels appear at $t=1$. Particles are shifted by one cell to the right at $t=2$. In particular, the particle ”d” appears in cell seven. The particle number distributions for indistinguishable and distinguishable particles are identical at $t=0$, 1 and 2. The pair annihilation reactions in cells 4, 5 and 6 are impossible during the third step because cells contain particles and antiparticles with different labels. For example, cell 4 contains particle ”a” and antiparticle ”b” at $t=2$. See the text for more details.
@@
Apparent teleportation within Cell Model is illustrated in Fig. 1. An example of the system’s time evolution with $V=9$ cells and periodic boundary conditions is shown. Cells contain ”background” particles that carry total energy, depicted by the grey band. Particles and antiparticles of interest are depicted by red and blue circles, respectively. For simplicity, we assume that the speed-of-light limit implies $\Delta=1$. The difference between the total number of particles and antiparticles is assumed to be conserved.
@@
The case of indistinguishable particles and antiparticles is presented in the plot a of Fig. 1. We postulate that only one particle exists at $t=0$, and it is in cell 3. Three pairs of particles and antiparticles appear in cells 4, 5 and 6 at $t=1$. At $t=2$, the particle number distribution is shifted by one cell to the right by the transport-local transition. This is allowed by the transport-locality inequalities (2.1). Then, at $t=3$, particle-antiparticle pairs annihilate, leaving a single particle in cell seven.
@@
The plot b of Fig. 1 shows an unsuccessful attempt to mimic the above process using distinguishable particles. In the example presented here, they are labelled by the letters ”a”, ”b”, ”c”, and ”d”. Similarly, antiparticles have unique labels. Only creating a particle and an antiparticle with identical labels is possible. Similarly, only a particle and an antiparticle with identical labels can annihilate.
@@
For the label-blind observer, the particle and antiparticle number distributions of indistinguishable and distinguishable particles are identical at $t=0,1$, and 2. But the similarity breaks at $t=3$. This is because annihilating particles and antiparticles with different labels is impossible. Of course, distinguishable particles can mimic the process depicted in the plot a if the transport-locality requirement is lifted. Particle ”a” can be ”teleported” by four cells in three steps from cell 333 to 777.
@@
The example construction of the process shown in Fig. 1 can be easily generalised, allowing for an arbitrary redistribution of indistinguishable boson-like particles and antiparticles in three time steps. This includes apparent teleportations - the processes which seemingly violate the speed of light limit.
@@
We note that the microscopic mechanism of the apparent teleportation illustrated in Fig. 1 a cannot be used for superluminal signalling. An observer in cell seven detecting the particle at $t=3$ cannot be sure it is caused by injecting a particle at $t=0$ in cell three. For example, this can occur when a particle and an antiparticle pair are created in cell six independently of events in cell three. Then a particle appears in cell seven.
@@
The probability of a given process depends on the dynamics which in the Cell Model is encapsulated within the transition matrix. One can speculate that at sufficiently high energy densities, the transition matrix of indistinguishable particles approaches maximum symmetry (the microstate symmetry Gazdzicki:2022zej ), and all processes have the same probability. This would imply that the system is “born in equilibrium” Hagedorn:1965st - the probability of any microstate appearing at any time is equal. This may explain the puzzle of fast equilibration in heavy-ion collisions at high energies Schlichting:2019abc . The apparent teleportation also may explain the quark-gluon paradox formulated in Ref. Miskowiec:2007qv . The transition of a large volume of quark-gluon plasma created in collisions of two atomic nuclei (see below for details) to colourless hadrons appears to require superluminal transport of colour charge to ensure local colour neutrality. Quarks and gluons are indistinguishable particles. Thus, the apparent teleportation may also serve as the explanation of the quark-gluon plasma paradox.
The above discussion motivates the question of whether the apparent teleportation, if it exists, can be observed experimentally. This question is addressed in the following subsection.
2.2 Observing apparent teleportation
@@
System of quarks and gluons created in nucleus-nucleus collisions at high energies is the highest energy density system ($\epsilon>1$ GeV/fm3 $\approx 1.6\cdot 10^{35}$ J/m3) created under conditions controlled in the laboratory Florkowski:Phenomenology . Its equilibrium state is called quark-gluon plasma (QGP) Shuryak:1980tp . The plasma expands and cools down, and the transition to colour-neutral hadrons - small bags of quarks and gluons -occurs at the hadronisation (transition) temperature $T_{\textrm{HAD}}\approx 150$ MeV Aoki:2006br .
@@
Six types of quarks (fermions) $u$, $d$, $s$, $c$, $t$ and $b$ and the corresponding antiquarks and the eight types of gluons (bosons) can be created during the collision. Massless gluons and light $u$ and $d$ quarks and antiquarks are the most abundant. Production of heavier quarks is suppressed even at the top collision energies of the CERN LHC ALICE:2013mez . The threshold for QGP creation in heavy-ion collisions is located at $\sqrt{s_{NN}}\approx 10$ GeV NA49:2002pzu ; NA49:2007stj . For a review, see also Refs. Gazdzicki:2010iv ; Andronov:2022cna . The fixed target experiments at the CERN SPS measure heavy-ion collisions in the energy range $\sqrt{s_{NN}}\approx 5-20$ GeV. At the top SPS energy, just above the QGP threshold energy, the mean number of light quarks and gluons produced in central collisions of two lead nuclei is on the order of 1000 Gazdzicki:1998vd . The corresponding numbers for strange and charm quarks are $\approx$ 100 Gazdzicki:1998vd and $\approx$ 1 Merzlaya:2024cbt , respectively.
@@
The local creation and annihilation of $c$ and $\bar{c}$ pairs implies that the energy and momentum scale of the interaction determines the $c$ and $\bar{c}$ space-time separation via the uncertainty principle. The scale ranges between:
@@
    The theoretical challenge: Quantum Chromodynamics, the theory of strong interactions, does not provide quantitative predictions for multi-particle correlations in heavy-ion collisions; for further discussion, see Ref. Gazdzicki:2023niq .
However, with the help of well-established heavy-ion models, it may be possible to provide experimental evidence of the apparent teleportation by measuring momentum correlations between $c$ and $\bar{c}$ quarks. The idea is rooted in our previous paper Gazdzicki:2023niq , which suggests the use of momentum correlations of charm-anticharm hadrons to extract information on spatial correlations of $c$ and $\bar{c}$ quarks at hadronisation. The paper reviews the experimental results on charm hadron correlations and collective flow. It also discusses the importance of having collisions with no more than one $c\bar{c}$ pair created. We note that the requirement of a single $c\bar{c}$ pair in a collision can be reconciled with the toy mechanism of the apparent teleportation presented in Sec. 2.1, assuming the coherent creation and annihilation of $c\bar{c}$ pairs in a short time interval.
[…]
@@
    Initial stage: a high-density quark-gluon plasma is created. Theoretical approaches addressing this stage are discussed in Ref. Schlichting:2019abc .
@@
    Expansion stage: the plasma expands Landau:1953gs , cooling down to the hadronisation temperature ($T_{\textrm{HAD}}\approx 150$ MeV in the local frame of the flowing matter). This stage is modelled using relativistic hydrodynamics Florkowski:Phenomenology .
@@
    Hadronisation stage: the plasma is converted into hadrons and resonances following statistical rules Hagedorn:1980kb ; Becattini:2005xt applied in the rest frame of a plasma fluid element. As a result, hadron momenta are determined by their flow acquired during the expansion stage and by the hadronisation (a local statistical process).
@@
[…]
Let us assume the $c$ and $\bar{c}$ quarks were created in the same space-time point within the collision light cone and moved to hadronisation as distinguishable particles with subluminal velocities. In this case, the creation point has to be located inside the overlap between the collision light cone and the past-light cones of $c$ and $\bar{c}$ quarks - the local-distinguishable (LD) condition. The $c\bar{c}$ pair depicted in Fig. 2 b fulfils the requirement, but the pair depicted in the plot c does not. The latter implies that the process was not local-distinguishable, and one needs the apparent teleportation to explain it.